# Materialist: Physically Based Editing Using Single-Image Inverse Rendering

### v1

@@ 3.2 Progressive Differentiable Rendering
$\mathcal{L}_{\text{re}}$ $=\mathcal{L}_{L1}(L_{o},I)+\mathcal{L}_{L2}(L_{o},I)$ (20)
$\mathcal{L}_{\text{cons}}$ $=\sum_{X\in\{\mathbf{A},\mathbf{R},\mathbf{M}\}}\mathcal{L}_{L1}(X_{o},X_{p})\,,$ (21)
@@
3.3 Opaque Material Editing
For $\mathbf{R,M,A}$ edits (excluding transparency), we use the SAM2 [47] model to segment the input image, creating a mask for the edit region. We can then modify values within this region and re-render an image. For $\mathbf{A}$ edits, we convert to HSV color space for intuitive color adjustments.
@@ 3.3 Opaque Material Editing
As described in Sec. 3.1, we reconstruct the mesh using the depth map, capturing only the visible front side of objects (see Fig. 3). This limits transparent object editing, as refractive effects cannot be fully modeled. While existing methods generate multi-view images for mesh reconstruction via NeRF [37, 31, 32, 54] or Gaussian Splatting [17, 23, 74], we aim to solve Material Transparency Editing directly within a single-view constraints, avoiding full scene reconstruction. We model transparent objects using DisneyBSDF [9] as:
$f_{\text{glass}}=\frac{\sqrt{\mathbf{A}_{\text{glass}}(\bm{x}_{s})}(1-F_{s})D_{s}G_{s}|\vec{h}\cdot\vec{\omega}_{o}||\vec{h}\cdot\vec{\omega}_{i}|}{|\vec{n}\cdot\vec{\omega}_{i}||\vec{n}\cdot\vec{\omega}_{o}|(\vec{h}\cdot\vec{\omega}_{i}+\eta\,\vec{h}\cdot\vec{\omega}_{o})^{2}}\,,$ (22)
@@
We use SAM2 [47] for object mask extraction and an inpainting model [48] to generate $\mathbf{A}_{\text{BG}}$. For an incoming ray of direction $-\vec{\omega}_{o}$ hitting the masked area, the direction of the transmitted ray $\vec{\omega}_{t}$ is computed using the law of refraction [22].
Assuming the object’s back normal is same as its front normal $\mathbf{N}_{p}(\bm{x}_{s})$, the world position after two refractions is
$\bm{x}_{2}=\bm{x}+d_{1}\,\vec{\omega}_{t1}+d_{2}\,\vec{\omega}_{t2}\,,$ (23)
@@
4 Experiments
This section evaluates the performance of our method.
4.1 Material Prediction Accuracy
@@
We evaluate the performance of our MatNet on the testset of the InteriorVerse [78] synthetic indoor dataset. As shown in Table 1, our method achieves better prediction accuracy across all evaluated metrics. Qualitative comparisons are shown in Fig. 4.
Table 1: Quantitative Evaluation of Material Prediction on Synthetic Dataset [78].
@@
Figure 4: Qualitative results of material prediction on synthetic dataset.
@@
Figure 5: Qualitative results of material prediction on the real world dataset IIW [4]. Zhu et al. [78] retained too many details from the original image while Kocsis et al. [24], due to their reliance on Stable Diffusion, altered image details excessively. Additionally, Kocsis et al.’s method struggles to handle specular highlights and facial features effectively.
Table 2: Quantitative evaluation, WHDR ($\delta=0.1$), of albedo predictions on the IIWild Dataset [4].
WHDR $\downarrow$
Li et al. [29] 0.342
Zhu et al. [78] 0.232
Kocsis et al. [24] 0.206
Ours 0.197
We assess the performance of albedo predictions using the IIW dataset [4], which includes sparse pairwise human annotations for albedo. For evaluation, we use the metric developed for the IIW dataset, Weighted Human Disagreement Rate (WHDR), which quantifies the error when predicted albedo values differ from human-provided annotations. As shown in Table 2, our model achieves better results with a lower WHDR score. Fig. 5 provides a qualitative comparison.
4.2 Environment Map Optimization
Currently, single-view envmap generation primarily relies on Stable Diffusion [44], which can only produce visually plausible envmaps. Optimization-based methods often require multi-view input [60, 72, 42] or use implicit lighting representations [69, 78, 24]. For optimization approaches, we selected DPI [35] as our benchmark, adapting it to single-view input and providing it with a mesh reconstructed using our method (Sec. 3.1).
Table 3: Quantitative evaluation of environment map optimization. Envmap measures the difference between the optimized and ground truth envmaps, while Rerender represents the difference between the rendered relit scene with the optimized envmap and the ground truth. SH ERROR calculates the error as the norm of the difference matrix between the spherical harmonics of the ground truth and optimized envmaps. The label mat indicates the material properties predicted by MatNet.
@@ 4.2 Environment Map Optimization
As our focus is not high-resolution envmaps, we limit the resolution to $32\times 16$, consistent with prior work [28, 12]. We evaluated three scenes, each with 60 envmap optimization results. Evaluation metrics included the difference between the optimized and the ground truth envmaps and the difference between relit scenes using the optimized and the ground truth envmaps. Both methods used ground truth material properties for relighting. As shown in Table 1, our method outperformed the others across all metrics, benefiting from the MatNet-predicted properties $\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}$ during envmap optimization. Qualitative results are in Fig. 6.
Figure 6: Qualitative results of envmap optimization.
@@
4.3 Object insertion and relighting
@@
Figure 7: Qualitative results of object insertion and relighting, compared with the multi-view method MAIR (Choi et al. [12]) and the single-view method by Li et al. [27]. Note that relight result2 has a realistic shadow.
@@
Figure 8: Comparison between ground truth (GT) and our object insertion result for real world image. The error in the reflective area mainly comes from the insufficient accuracy of optimized envmap, as explained in Sec. 4.2.
4.4 Material Editing
For transparency editing evaluation, we used real-world photo comparisons. A 3D-printed opaque jug, identical in size to the scanned mesh from Sec. 4.3, was placed alongside a glass jug at a fixed camera position, and two photos were captured as ground truth for transparency editing. Since Alchemist [53], closely related to our work, is not open-sourced, we compared our method with iPix2Pix [7] and Stable Diffusion inpainting [48]. For iPix2Pix, text prompts were used for editing, while Stable Diffusion inpainting received both a mask of the jug region and a text prompt. Results are shown in Fig. 9.
@@ 4.4 Material Editing
Figure 9: Material transparency editing compared with iPix2Pix [7] and Stable Diffusion inpainting [48]. Our method can more accurately replicate the distortion effects caused by light refraction.
Figure 10: Qualitative evaluation of material editing. Our method enables albedo editing to various colors, whereas Alchemist [53] is limited to white. For roughness editing, our approach follows physical principles—when roughness is set to 0.2, a realistic reflection of the stone appears at the bottom of the bottle. Our method preserves Charizard’s color during metallic edits and shows that increasing metallic alone darkens the material. In transparency editing, our approach accurately captures light refraction and maintains the vase’s appearance when transparency is increased.
@@
For albedo, roughness, and metallic evaluation, since Alchemist [53] is not open-sourced, we extracted images from its paper and applied our material editing for comparison. These images are notably out-of-domain for indoor scenes. While MatNet predicts accurately for indoor scenes, only albedo predictions remain reliable, with roughness and metallic predictions being less consistent. However, differential rendering allows us to perform inverse rendering on real photos using only albedo. To compensate for unreliable roughness and metallic predictions, we manually set priors for each image to improve optimization. Results are shown in Fig. 10. Detailed settings and user study are provided in the supplementary material.
4.5 Human Face Relighting
@@
We also tested the performance of our method on human face relighting to demonstrate its potential. We used Material Net, trained on the indoor dataset and not specifically optimized for the human face dataset, to predict the $\mathbf{A},\mathbf{R},\mathbf{M}$, followed by progressive optimization. The optimized results were used to relit. It is important to note that, despite the seemingly good results (as shown in Figure 11), we did not model the subsurface scattering, which is critical for realistic face rendering. In our approach, subsurface scattering component is fixed within the albedo during optimization and does not change with lighting, thus making the rendered results inaccurate. This could be possibly improved in future work by integrating subsurface scattering into the rendering pipeline. Additionally, if there are specular highlights on the face, they too would be fixed in the albedo during the optimization process.
Figure 11: Our method was applied to relight real human faces (FFHQ). When the face lacks strong highlights, the relighting results are more accurate (above). However, in cases where highlights are present (below), these highlights tend to be baked into the albedo during optimization, causing them to persist even after the lighting conditions are altered.
5 Ablations
@@
We tested the results of envmap optimization under two conditions: providing MatNet predictions versus setting $\mathbf{A}=\mathbf{R}=0.5,\mathbf{M}=0.1$. As shown in Table 3, providing MatNet predictions significantly improves the accuracy of envmap optimization.
@@
We also conducted an ablation study on the value of $\delta$ in Eq. 19. Using 20 randomly selected images from the IIW dataset [4], we tested different $\delta$ values and compared the rerendered images with the input images. We evaluated two strategies: simultaneous optimization of $\mathbf{ARM}$ and sequential optimization ($\mathbf{RM\&A}$), where $\mathbf{RM}$ is optimized first, followed by $\mathbf{A}$. To save time, we applied early stopping, halting optimization if $\mathcal{L}_{\text{re}}$ reduced by less than 5% over 20 consecutive steps. Results are shown in Table 4. For additional studies, see the supplementary material.
Table 4: Ablation of $\delta$ in Eq. 19. Comparing the rerender error.
@@
6 Limitations and Discussion
Figure 12: Super-resolution can improve image quality while preserving overall structure, such as lighting and shadows. However, Stable Diffusion-based super-resolution may alter some details, as seen in the excessive modifications to the statue in the right image.
Since our images are rendered and mesh reconstruction limitations can cause artifacts along object edges, using super-resolution (SR) [48, 59, 65, 66] may help reduce these artifacts. However, SR models often overly modify the image, as shown in Fig. 12. Improving SR models by giving additional input could potentially mitigate this issue.
Finally, physically based in this paper refers to physically based rendering (PBR), a computer graphics approach that may not strictly follow physical laws. For more general limitations, please refer to the supplementary material.
7 Conclusion
In this work, we present a novel pipeline for physically-based inverse rendering that combines neural network predictions with differentiable rendering to optimize material properties and lighting conditions from a single image. Our approach enables accurate relighting, transparent object insertion, and material editing tasks, including transparency editing without requiring complete geometry.
Supplementary Material
8 More Ablations
@@
8.1 Optimize SH coefficients instead of envmap

### v2

@@ 3.2. Progressive Differentiable Rendering
(13) $\mathcal{L}_{\text{re}}$ $=\mathcal{L}_{L1}(L_{o},I)+\mathcal{L}_{L2}(L_{o},I)$
(14) $\mathcal{L}_{\text{cons}}$ $=\sum_{X\in\{\mathbf{A},\mathbf{R},\mathbf{M}\}}\mathcal{L}_{L1}(X_{o},X_{p})\,,$
@@
3.3. Opaque Material Editing
For $\mathbf{R,M,A}$ edits (excluding transparency), we use the SAM2 (Ravi et al., 2024) model to segment the input image, creating a mask for the edit region. We can then modify values within this region and re-render an image. For $\mathbf{A}$ edits, we convert to HSV color space for intuitive color adjustments.
@@ 3.3. Opaque Material Editing
As described in Sec. 3.1, we reconstruct the mesh using the depth map, capturing only the visible front side of objects (see Fig. 3). This limits transparent object editing, as refractive effects cannot be fully modeled. While existing methods generate multi-view images for mesh reconstruction via NeRF (Mildenhall et al., 2021; Liu et al., 2024; Long et al., 2024; Shi et al., 2023) or Gaussian Splatting (Guédon and Lepetit, 2024; Kerbl et al., 2023; Yu et al., 2024), we aim to solve Material Transparency Editing directly within a single-view constraints, avoiding full scene reconstruction. We model transparent objects using DisneyBSDF (Burley, 2015) as:
(15) $f_{\text{glass}}=\frac{\sqrt{\mathbf{A}_{\text{glass}}(\bm{x}_{s})}(1-F_{s})D_{s}G_{s}|\vec{h}\cdot\vec{\omega}_{o}||\vec{h}\cdot\vec{\omega}_{i}|}{|\vec{n}\cdot\vec{\omega}_{i}||\vec{n}\cdot\vec{\omega}_{o}|(\vec{h}\cdot\vec{\omega}_{i}+\eta\,\vec{h}\cdot\vec{\omega}_{o})^{2}}\,,$
@@
We use SAM2 (Ravi et al., 2024) for object mask extraction and an inpainting model (Rombach et al., 2022) to generate $\mathbf{A}_{\text{BG}}$. For an incoming ray of direction $-\vec{\omega}_{o}$ hitting the masked area, the direction of the transmitted ray $\vec{\omega}_{t}$ is computed using the law of refraction (Kay and Greenberg, 1979). Assuming the object’s back normal is same as its front normal $\mathbf{N}_{p}(\bm{x}_{s})$, the world position after two refractions is
(16) $\bm{x}_{2}=\bm{x}+d_{1}\,\vec{\omega}_{t1}+d_{2}\,\vec{\omega}_{t2}\,,$
@@
4. Experiments
It is important to note that MatNet serves as a support component within our pipeline; we do not directly compete with image decomposition works. Rather, more precise material predictions enhance the optimization of envmaps through the utilization of our pipeline.
4.1. Material Prediction Accuracy
@@
We evaluate the performance of our MatNet on the testset of the InteriorVerse (Zhu et al., 2022b) synthetic indoor dataset. As shown in Table 1, our method achieves better prediction accuracy across all evaluated metrics. Qualitative comparisons are shown in Fig. 13.
Table 1. Quantitative Evaluation of Material Prediction on Synthetic Dataset (Zhu et al., 2022b).
@@
We evaluate albedo predictions on the IIW dataset (Bell et al., 2014) using the Weighted Human Disagreement Rate (WHDR), which measures errors against human annotations. As shown in Table 2, our model achieves a lower WHDR score and the fastest inference speed. Qualitative results are provided in Fig. 4.
Figure 4. Qualitative results on the IIW dataset (Bell et al., 2014) highlight key differences: Zhu et al. (Zhu et al., 2022b) retain excessive details, Kocsis et al. (Kocsis et al., 2024) distort textures and struggle with specular highlights and facial features, while Zeng et al (Zeng et al., 2024). over-smooth, losing critical details like facial features and background information. These issues underscore the challenge of balancing detail preservation and realistic material prediction.
Table 2. Quantitative evaluation of albedo predictions on the IIWild Dataset (Bell et al., 2014) using WHDR ($\delta=0.1$). Time denotes the average inference time in seconds per image on one H100 (80G), with parentheses indicating the speedup factor relative to the slowest method. Notably, Zeng et al. (Zeng et al., 2024) achieve a better WHDR score by over-smoothing images, but this comes at the cost of losing significant details, as shown in Fig. 4.
WHDR $\downarrow$ Time $\downarrow$
Li et al. (Li et al., 2022) 0.342 0.221 (35.7$\times$)
Zhu et al. (Zhu et al., 2022b) 0.232 0.229 (34.4$\times$)
Zeng et al. (Zeng et al., 2024) 0.190 7.880 (1.0$\times$)
Kocsis et al. (Kocsis et al., 2024) 0.206 6.730 (1.2$\times$)
Ours 0.197 0.204 (38.7$\times$)
4.2. Environment Map Optimization
Currently, single-view envmap generation primarily relies on DM (Phongthawee et al., 2024), which can only produce visually plausible envmaps. Optimization-based methods often require multi-view input (Verbin et al., 2024; Yu et al., 2023a; Park et al., 2020) or use implicit lighting representations (Yao et al., 2022; Zhu et al., 2022b; Kocsis et al., 2024). For optimization approaches, we selected DPI (Lyu et al., 2023) as our benchmark, adapting it to single-view input and providing it with a mesh reconstructed using our method (Sec. 3.1).
Table 3. Quantitative evaluation of envmap optimization. Envmap measures the difference between the optimized and ground truth envmaps, while Rerender represents the difference between the rendered relit scene with the optimized envmap and the ground truth. SH ERROR calculates the error as the norm of the difference matrix between the spherical harmonics of the ground truth and optimized envmaps. The label mat indicates the material properties predicted by MatNet.
@@ 4.2. Environment Map Optimization
Figure 5. Qualitative results of envmap optimization.
As our focus is not high-resolution envmaps, we limit the resolution to $32\times 16$, consistent with prior work (Li et al., 2021; Choi et al., 2023). We evaluated three scenes, each with 60 envmap optimization results. Evaluation metrics included the difference between the optimized and the ground truth envmaps and the difference between relit scenes using the optimized and the ground truth envmaps. Both methods used ground truth material properties for relighting. As shown in Table 1, our method outperformed the others across all metrics, benefiting from the MatNet-predicted properties $\mathbf{A}_{p},\mathbf{R}_{p},\mathbf{M}_{p}$ during envmap optimization. Qualitative results are in Fig. 5.
@@
4.3. Object insertion and relighting
@@
Figure 6. Qualitative results of object insertion and relighting, compared with the multi-view method MAIR (Choi et al. (Choi et al., 2023)) and the single-view method by Li et al. (Li et al., 2020). Note that relight result2 has a realistic shadow.
@@
Figure 7. Material transparency editing compared with iPix2Pix (Brooks et al., 2023) and DM inpainting (Rombach et al., 2022). Transmission is set to 0.8, and object color is set to white. Our method can more accurately replicate the distortion effects caused by light refraction.
Figure 8. Comparison between ground truth (GT) and our object insertion result for real world image. The error in the reflective area mainly comes from the insufficient accuracy of optimized envmap, as explained in Sec. 4.2.
4.4. Shadow Effects
Recent advancements in diffusion model-based neural rendering, exemplified by the state-of-the-art work DiffusionRenderer (Liang et al., 2025), have demonstrated significant improvements in shadow quality. However, due to its reliance on a geometry-free approach, our method, which leverages 3D geometry, continues to achieve superior shadow accuracy, as shown in Fig 15.
4.5. Material Editing
For transparency editing evaluation, we used real-world photo comparisons. A 3D-printed opaque jug, identical in size to the scanned mesh from Sec. 4.3, was placed alongside a glass jug at a fixed camera position, and two photos were captured as ground truth for transparency editing. Since Alchemist (Sharma et al., 2024), closely related to our work, is not open-sourced, we compared our method with iPix2Pix (Brooks et al., 2023) and DM inpainting (Rombach et al., 2022). For iPix2Pix, text prompts were used for editing, while DM inpainting received both a mask of the jug region and a text prompt. Results are shown in Fig. 7.
@@ 4.5. Material Editing
For albedo, roughness, and metallic evaluation, since Alchemist (Sharma et al., 2024) is not open-sourced, we extracted images from its paper and applied our material editing for comparison. These images are notably out-of-domain for indoor scenes. While MatNet predicts accurately for indoor scenes, only albedo predictions remain reliable, with roughness and metallic predictions being less consistent. However, differential rendering allows us to perform inverse rendering on real photos using only albedo. To compensate for unreliable roughness and metallic predictions, we manually set priors for each image to improve optimization. Results are shown in Fig. 15. Detailed settings and user study are provided in the supplementary material.
4.6. Physical Simulation
In contrast to neural rendering approaches (Liang et al., 2025), our method, benefiting from complete geometry, enables unique physically-based simulation capabilities. Fig. 13 illustrates a simulation of virtual objects falling in varying illumination conditions, reconstructed from a single image inverse rendering using our methods.
4.7. Human Face Relighting
@@
We evaluated our method on human face relighting to showcase its potential. Using Material Net, trained on an indoor dataset and not fine-tuned for faces, we predicted $\mathbf{A},\mathbf{R},\mathbf{M}$ and performed progressive optimization for relighting. While results appear promising (Figure 13), our approach does not model subsurface scattering, which is crucial for realistic face rendering. In our method, subsurface scattering is fixed within the albedo during optimization and remains static under lighting changes, limiting accuracy. Similarly, specular highlights on the face are also fixed in the albedo. Future work could improve realism by integrating subsurface scattering into the rendering pipeline.
5. Ablations
@@
We tested the results of envmap optimization under two conditions: providing MatNet predictions versus setting $\mathbf{A}=\mathbf{R}=0.5,\mathbf{M}=0.1$. As shown in Table 3, providing MatNet predictions significantly improves the accuracy of envmap estimation.
@@
We also conducted an ablation study on the value of $\delta$ in Eq. 12. Using 20 randomly selected images from the IIW dataset (Bell et al., 2014), we tested different $\delta$ values and compared the rerendered images with the input images. We evaluated two strategies: simultaneous optimization of $\mathbf{ARM}$ and sequential optimization ($\mathbf{RM\&A}$), where $\mathbf{RM}$ is optimized first, followed by $\mathbf{A}$. To save time, we applied early stopping, halting optimization if $\mathcal{L}_{\text{re}}$ reduced by less than 5% over 20 consecutive steps. Results are shown in Table 4. For additional studies, see the supplementary material.
Furthermore, while lower-resolution envmaps typically exert a limited influence on rendering results, this is not the case for highly reflective surfaces. We conducted an ablation study to assess the impact of varying envmap resolutions on such surfaces. Importantly, we maintained ground truth material properties and optimized solely with respect to the envmap. The results is shown in Fig. 13.
Table 4. Ablation of $\delta$ in Eq. 12. Comparing the rerender error.
@@
6. Limitations and Conclusion
Since our images are rendered and mesh reconstruction limitations can cause artifacts along object edges, using super-resolution (SR) (Rombach et al., 2022; Team, 2024; Wang et al., 2021b; Wu et al., 2024) may help reduce these artifacts. However, SR models often overly modify the image, as shown in Fig. 13. Improving SR models by giving additional input could potentially mitigate this issue. For more general limitations, please refer to the supplementary material.
We present a novel method for inverse rendering that combines neural network predictions with differentiable rendering to optimize material properties and lighting conditions from a single image. Our approach enables more accurate envmap estimation, scene level relighting, object insertion, and material editing including transparency editing without requiring complete geometry.
Acknowledgements.
This project has received funding from the European Union’s Horizon 2020 research and innovation programme under the Marie Skłodowska-Curie grant agreement No. 956585 (PRIME).
This project has received funding from Innovation Fund Denmark, project 0223-00041B (ExCheQuER).
Supplementary Materials
Appendix A Limitations
In this section, we discuss potential limitations of our proposed method from a broader, user-oriented perspective. It is important to note that some points, such as the need for per-image optimization, are not typically viewed as limitations within the differentiable rendering field. However, from a user’s viewpoint, this may make our approach less convenient than DM-based methods (Sharma et al., 2024; Brooks et al., 2023).
1. As with all differentiable rendering methods, our approach requires optimization for each image to obtain the lighting and optimized material properties, which might be quite time-consuming. Depending on MatNet’s prediction accuracy, optimization takes between 5 to 30 minutes. For out-of-domain images (i.e., non-indoor scenes) or higher-quality requirements, optimization time may increase. This limits our method’s ability to achieve the rapid, batch image editing enabled by stable diffusion-based techniques (Brooks et al., 2023). Training MatNet on a broader dataset for more accurate predictions could help mitigate this limitation.
4. For Material Transparency Editing, complex object geometry can lead to inaccurate refraction distortions, as our method assumes only two refractions. Single-view mesh reconstruction (Liu et al., 2024; Long et al., 2024; Shi et al., 2023) may yield more accurate refractions for complex shapes. Besides, since we simulate refraction without the rays truly passing through the object, shadows cast by transparent objects may lack accuracy, and does not have caustic effect. (Note that this problem does not exist for object insertion tasks with complete geometry.)
6. A key limitation of our method, as with all single-image Differentiable Rendering approaches, is the inherent ambiguity in disentangling object colors, which can arise from either lighting or material properties. While our optimization process operates under these constraints, we mitigate this ambiguity through the use of priors and techniques (Sec. E). However, fully resolving this challenge remains beyond the scope of our work. Despite this limitation, our method achieves superior image editing results, as demonstrated in Figures. In an ideal scenario, if MatNet could infer ground truth materials directly, optimization would be unnecessary. Nonetheless, our approach strikes a practical balance between performance and feasibility under these inherent constraints.
Figure 16. Super-resolution often struggles with refraction effects. As shown, both Super Resolution 1 and Super Resolution 2 exhibit noticeable refraction issues, requiring multiple runs to achieve a satisfactory result, as seen in Super Resolution 3.
Appendix B More Ablations
@@
B.1. Optimize SH coefficients instead of envmap